# Measuring the Carbon Footprint of Cryptographic Privacy-Enhancing Technologies

### v1

@@ Abstract
Privacy-enhancing technologies (PETs) have attracted significant attention in response to privacy regulations, driving the development of applications that prioritize user data protection. At the same time, the information and communication technology (ICT) sector faces growing pressure to reduce its environmental footprint, particularly its carbon emissions. While numerous studies have assessed the energy footprint of various ICT applications, the environmental footprint of cryptographic PETs remains largely unexplored.
Our work addresses this gap by proposing a standardized methodology for evaluating the carbon footprint of PETs. To demonstrate this methodology, we focus on PETs supporting client-server applications as they are the simplest to deploy. In particular, we measure the energy consumption and carbon footprint increase induced by five cryptographic PETs (compared to their non-private equivalent): HTTPS web browsing, encrypted machine learning (ML) inference, encrypted ML training, encrypted databases, and encrypted emails. Our findings reveal significant variability in carbon footprint increases, ranging from a twofold increase in HTTPS web browsing to a 100,000-fold increase in encrypted ML.
Our study provides essential data to help decision-makers assess privacy-carbon trade-offs in such applications. Finally, we outline key research directions for developing PETs that balance strong privacy protection with environmental sustainability.
Keywords:
Privacy-Enhancing Technologies Carbon Footprint CryptographyEncrypted ML Encrypted Databases HTTPS.
@@ 1 Introduction
On the other hand, various ML paradigms have been studied under the lens of carbon footprint, including differentially private ML [35, 38] and Federated ML [19, 40]; two paradigms related to privacy-preserving ML. Our work is complementary to these research works because we notably measure the footprint of cryptographic techniques applied to privacy-preserving ML.
@@
We present a standard methodology to analyze the footprint of a PET. In particular, we detail a taxonomy of the possible overheads induced by a PET. This methodology enables to answer two key questions: (1) What is the carbon footprint of a given cryptographic PET? (2) What is the relative increase compared to a non-private equivalent?
To demonstrate this methodology, we measure the energy consumption and carbon footprint of five cryptographic PETs: HTTPS, encrypted ML inference, encrypted ML training, encrypted databases, and encrypted emails. We focus on PETs supporting client-server applications, because they are the easiest to deploy. In particular, we highlight the relative carbon footprint difference between the PET and its non-private equivalent (i.e., the carbon footprint of the privacy enhancement). We show highly variable carbon footprint increases from a $\times 2$ increase in HTTPS to a $\times 100,000$ in encrypted ML.
Finally, we discuss promising research directions to build sustainable and privacy-enhancing technologies.
Our goal is not to categorize PETs between environmentally acceptable and unacceptable. To fulfil carbon emission goals, it is not mandatory to reduce the carbon emissions of all technologies: some highly energy-consuming technologies can be considered “worth the emissions” because their service is essential. Our work simply provides orders of magnitude enabling decision makers to assess the privacy-carbon trade-offs of specific services. Such figures have become essential in the public debate as press articles rely on such estimations and measurements to provide high-level perspectives on sustainability issues [18].
2 Measuring the carbon footprint of a PET
@@
Taking inspiration from Schmidt et al. [41] who introduced carbon footprint measurement to the ML community, this section applies the LCA methodology to PETs and present simple analysis tools for PET researchers.
2.1 Background: Life-Cycle Assessment
@@
While LCA focuses on physical objects, it is possible to extend this approach to digital services. However, LCA on such services focuses mostly on the service usage. For example, for an ML service, an LCA studies the impacts of the model training [23, 30, 41], and of the inference [14].
@@
Like most works on the carbon footprint of ICTs [17, 23, 33, 35, 36, 38, 40, 41], our LCA focuses on the service usage. Therefore, our experiments quantify the impact in terms of carbon footprint and energy consumption.
2.2 Taxonomy of carbon footprint overhead in PET
Our goal is to measure the “footprint overhead” induced by PET usage; i.e., the footprint increase between a PET and a non-private equivalent. This footprint increase can have several origins in PETs. To compare the carbon footprint of PETs, we distinguish four types of overhead:
  • Computational overhead: any additional computation (e.g., encryption).
  • Communication overhead: any additional communication (e.g., if the ciphertext is larger than the plaintext).
  • Infrastructure overhead: any server duplication (e.g., like the existence of two non-colluding servers in privacy-preserving telemetry [9]).
  • Hardware overhead: any PET-specific hardware requirement (e.g., TEE).
2.3 Our scope: client-server applications
As there exists plenty of PETs, we demonstrate our methodology on a specific subset and leave the rest of them for future works. Our work focuses on PETs in a client-server setup because they are the easiest to deploy. This simplicity makes such PETs attractive to companies providing “Software as a Service” because they simply need to update their code and do not require major infrastructure changes. For example, encrypted databases (such as MongoDB Queryable Encryption) fit this category because this PET only requires updating the software.
Figure 1: Our scope: non-private vs. privacy-enhanced client-server technology
Thus, we compare the computation costs of a client-server PET deployment to the non-private equivalent. More precisely, we highlight the relative footprint increase due to the privacy enhancement. This focus on client-server setups implies that the studied PETs induce no infrastructure overhead because, like the non-private equivalent, such PETs only require a client and a server.
Moreover, we exclude communication overheads from our scope because no existing work enables quantifying appropriately this overhead. Indeed, measuring the footprint of network communications is still an open research problem. No work has yet estimated the complete footprint of a client-server communications because it requires computing the amortized energy consumption of all network devices (e.g., routers, switches, antennas, wires, etc.) used to transmit a bit of information. Ficher et al. [15] started exploring this question, but they studied only the communication between two servers in the same network backbone. Further research is still necessary to obtain a general result.
Anyway, client-server PETs usually induce minor communication overheads. For example, in HTTPS web browsing, the size of the ciphertext sent to the server is comparable in size to the plaintext (plus a one-time secret-key exchange). Their communication overhead is anecdotal compared to communication overheads in “communication-intensive PETs” such as multiparty computations (MPC).
Finally, we leave for future works PETs based on cryptographic hardware (e.g., Trusted Execution Environments [39]). Their need for a dedicated hardware calls for a fine-grained analysis of this “hardware overhead.”
In summary, due to the properties of client-server PETs, our work only need to focus on the computational overhead. This scope already fits many PETs and our experiments covers 5 of them. Future works would be necessary to assess the carbon footprint of all possible PETs. Our research provides a clear framework supporting such future research and demonstrate its practicality on the PETs analyzed in Section 3.
@@ 2.4 Software-based measurement
In addition to energy consumption, our scope also expects carbon emission measurement. To compute the carbon emissions, one needs to multiply the energy consumption to the “carbon intensity” of the country in which the server/client is located. The carbon intensity corresponds to the amount of greenhouse gas emitted per kWh of electricity produced. This information is directly available in some public databases [13]. In other words, the energy consumption and the energy emissions are perfectly correlated. This carbon emission estimation is commonly used in ML research [23, 41].
2.5 Is the runtime a good proxy for the carbon footprint?
As research papers typically benchmark the runtime of their algorithms, one may want to extrapolate the carbon footprint based on such measurements. While runtime is correlated to energy consumption, other parameters influence it. Thus, we cannot trivially deduce the energy consumption based on the runtime.
@@ 2.5 Is the runtime a good proxy for the carbon footprint?
This observation shows that specialized energy measurement techniques (such as RAPL [26]) are essential to quantify the carbon footprint and cannot be obtained using simpler methods such as runtime measurement. For example, RAPL-based measurements accounts for the energy consumption of all cores.
In conclusion, we provide a complementary perspective to existing PET papers because we report carbon and energy measurements that cannot be precisely extrapolated from existing experimental results. Our work echoes with recent ML works [23] that promoted a systematic reporting of carbon footprint and energy consumption in ML papers (in addition to existing runtime measurements).
@@ 2.6 Experimental setup
As the hardware can slightly change the energy consumption of an algorithm, one may wonder whether we should execute our experiment on different devices. However, remember that we mainly want to measure the relative footprint difference between private and non-private deployments. A standard hardware change would influence the absolute values, but the relative difference should not be significantly changed. This LCA on a single hardware platform aligns with common practices in studies performing LCA on ML algorithms [38, 19].
@@
We use the software CodeCarbon [41] to measure the energy consumption and infer the subsequent carbon emissions. This software uses a RAPL-based measurement, and measures RAM, CPU, and GPU consumption.
We configured CodeCarbon to poll the energy consumption every 1 ms (like in [24]). Moreover, we chose the Netherlands as reference country to compute the carbon emissions. The Netherlands have a quite “average carbon intensity” within Europe (i.e., 268 g eq. CO2 per kWh in 2023) compared to the low carbon intensity of France (56 g eq. CO2 per kWh) and the high carbon intensity of Poland (662 g eq. CO2 per kWh). However, the choice of a country does not matter too much: we want to measure the relative increase. Reproducing the experiment in another country would change the absolute carbon emission but not the relative difference.
As explained above and illustrated in Figure 2, the energy and carbon emissions are then perfectly correlated in our experiments. Thus, some figures (as Figures 4(a) and 5(a)) only report the energy consumption in order to comply with space limitations. The reader can easily extrapolate the carbon emissions of a PET in any country by multiplying the energy consumption (reported in our figures) with the carbon intensity of the desired country [13].
3 Experimental results
This section measures the relative difference between the footprint of different PETs and the footprint of their non-private equivalent. Our measurement aggregates the computational costs of the client and the server.
Our experiments systematically use the most default configuration proposed in the documentation of the tested software. Thus, the measurements reported in our paper correspond to the expected default behavior. For implementation details, we refer to our publicly available codebase: https://github.com/MarcT0K/privacy-carbon-experiments
3.1 Encrypted Machine Learning inference
With the increasing use of ML, there is a growing concern regarding the use of personal data in these systems. To address this problem, many works [12, 34] have proposed privacy-preserving ML protocols to support these applications while preserving data privacy. Among all ML operations, ML inference (i.e., evaluating data on a trained ML model) is a particularly important operation for practitioners. Indeed, many companies have trained a powerful model (e.g., a Large Language Model) and provide “Inference as a Service”: customers send input data and the company returns the model output.
However, traditional “Inference as a Service” requires sharing personal data in plaintext with the service provider. To enhance privacy, researchers have designed encrypted ML frameworks [8] (based on Fully Homomorphic Encryption) to perform ML operations (in particular inference) under encryption.
Software libraries. Our goal is to measure the energy consumption of an encrypted inference compared to a classic plaintext inference. We use the software library Concrete ML developed by the company Zama. This library is stable and promoted by the company for real-world use cases. We compare the encrypted inference provided by Concrete ML to the plaintext ML inference performed by Scikit Learn (a popular ML framework). Note that Scikit Learn is also used as baseline by Zama in their accuracy benchmarks.
@@ 3.1 Encrypted Machine Learning inference
Figure 2: Average energy consumption and carbon emissions of encrypted and plaintext ML inference using various classification models (100 samples, 30 features).
Figure 3: Average energy consumption of encrypted and plaintext ML inference using various regression models (100 samples, 30 features).
Results. Figures 2 and 3 presents measurements for various classification and regression models. These results are average measurements over 100 samples; each sample having 30 features. On linear models (e.g., Logistic regression), the encrypted inference is $100\times$ more expensive than the plaintext inference. On tree models (Decision Tree, Random Forest, and XGB), the encrypted inference is at least $100,000\times$ more expensive than the plaintext inference. Encrypted inference on neural networks is also $100,000\times$ more expensive than plaintext.
@@
3.2 Encrypted Machine Learning training
Next to inference, ML training is the other operation attracting a lot of privacy concerns. Like for inference, researchers have designed frameworks [12, 16] to execute these algorithms directly on encrypted data. These frameworks usually rely on the same cryptographic primitives as encrypted inference.
Software libraries. Even though the Concrete ML library focuses on encrypted inference, it also implements a few training algorithms. In particular, it enables training a logistic regression under encryption. We then compare this encrypted training to the logistic regression of Scikit Learn.
Data. Like for inference, we use Scikit Learn to generate synthetic data with a varying number of samples and features.
@@ 3.2 Encrypted Machine Learning training
Results. Figure 5(a) compares the encrypted and plaintext training for a varying number of features. The encrypted training is at least $100,000\times$ more expensive and its costs increase faster with the number of features.
Contrary to inference, the logistic regression training is not influenced by the number of samples. Training algorithms are iterative and process a fixed-sized batch during each iteration. Thus, their cost only depends on the batch size and number of iterations. Figure 5(b) confirms this behavior experimentally.
Discussions. Like secure ML inference, secure training induces massive footprint increase. Similarly, there is room for optimizations.
However, other (non-cryptographic) approaches to privacy-preserving ML training exist. For example, de Reus et al. [38] studied the carbon footprint of synthetic data generation in ML context. Synthetic data generation [25] is a pre-processing step used in privacy-preserving statistics and ML: instead of encrypting its private data, the data owner generates synthetic data based on its data and can transmit the synthetic data in plaintext. In this paradigm, the server can then process plaintext data without privacy issues as it has only access to synthetic data (and not the initial private data). Some techniques such as [25] ensure that the synthetic data does not leak private information.
@@
This comparison emphasizes that the PET choice has a major impact on the carbon footprint. It also highlights an interesting future work: comparing all privacy-preserving ML paradigms to identify the most energy-efficient one.
3.3 Encrypted databases
@@
Gui et al. [20] already compared the runtime of SWiSSSE to Redis. We reproduce their benchmark using new metrics: energy and carbon.
@@
Results. Figure 6 shows the results of this benchmark for varying database sizes. We observe that the footprint of the encrypted database is nearly ten times higher than the footprint of the plaintext database. Moreover, the footprint of the encrypted database increases slightly faster in function of the database size.
Figure 6: Average energy consumption and carbon emissions of encrypted and plaintext queries (SWiSSSE vs. Redis) for varying database sizes (1000 queries).

### v2

# Energy Consumption of TLS, Searchable Encryption and Fully Homomorphic Encryption

@@ Abstract
Privacy-enhancing technologies (PETs) have attracted significant attention in response to privacy regulations, driving the development of applications that prioritize user data protection. At the same time, the information and communication technology (ICT) sector faces growing pressure to reduce its environmental footprint, particularly its energy consumption. While numerous studies have assessed the energy consumption of ICT applications, the environmental impact of cryptographic PETs remains largely unexplored.
This work investigates this question by measuring the energy consumption increase induced by three PETs compared to their non-private counterparts: TLS, Searchable Encryption, and Fully Homomorphic Encryption (FHE). These technologies were chosen for two reasons. First, they cover different maturity levels –from the widely deployed TLS protocol to the emerging FHE schemes– allowing us to examine the influence of maturity on energy consumption. Second, they each have well-established applications in industry: web browsing, encrypted databases, and privacy-preserving machine learning.
Our results reveal highly variable energy consumption increases, ranging from $2\times$ for TLS to $10\times$ for Searchable Encryption and $100{,}000\times$ for FHE. Our experiments demonstrate a simple and reproducible methodology, based on existing open-source software, to quantify the energy costs of PETs. They also highlight the wide spectrum of energy demands across technologies, underscoring the importance of further research on sustainable PET design. Finally, we discuss orthogonal research directions, such as hardware acceleration, to outline promising directions toward sustainable PETs.
@@ 1 Introduction
On the other hand, various ML paradigms have been studied under the lens of energy consumption, including differentially private ML [30, 32] and Federated ML [15, 34]; two paradigms related to privacy-preserving ML. Our work is complementary to these research works because we notably measure the consumption of cryptographic techniques applied to privacy-preserving ML.
@@
We present a simple and reproducible methodology (used notably in ML) to measure the energy consumption of PETs. Using this methodology, we measure the energy consumption of three PETs: TLS, Searchable Encryption, and Fully Homomorphic Encryption. These three PETs were selected because they combine practical relevance with different maturity levels. On the one hand, they have well-identified industrial applications: web browsing, encrypted databases, and privacy-preserving ML. On the other hand, they have different maturity levels from the widely deployed TLS protocol to the emerging FHE schemes. Our experiments highlight the relative increase between the PET and its non-private equivalent. We observe increases ranging from a twofold overhead with TLS to a 100,000-fold overhead with FHE. Finally, we discuss several orthogonal questions (including hardware acceleration) to identify promising directions for sustainable PET design.
Our goal is not to categorize PETs between environmentally acceptable and unacceptable. To fulfil carbon emission goals, it is not mandatory to reduce the carbon emissions of all technologies: some highly energy-consuming technologies can be considered “worth the emissions” because their service is essential. Our work simply provides orders of magnitude enabling decision makers to assess the privacy-efficiency trade-offs of specific services. Such figures have become essential in the public debate as press articles rely on estimations and measurements from the scientific literature to provide high-level perspectives on sustainability issues [14].
2 Methodology
Our work adopts a methodology commonly used in scientific studies (notably in ML research [32, 35]) to measure the energy consumption of various algorithms. This section presents the methodology, its underlying assumptions, and the general philosophy guiding its use in our study.
2.1 Life-Cycle Assessment
@@
While LCA initially focused on physical objects, it is possible to extend this approach to digital services. However, LCA on such services focuses mostly on the service usage. For example, for an ML service, an LCA studies the impacts of the model training [19, 26, 35] and inference [11].
@@
Like most related works on the environmental footprint of ICTs [13, 19, 29, 30] [31, 32, 34, 35], our LCA focuses on the service usage (especially the energy consumption). Therefore, our experiments quantify the environmental impact in terms of energy consumption.
@@ 2.2 Software-based measurement
From the energy consumption, note that it is possible to estimate the carbon emission. To compute the carbon emissions, one needs to multiply the energy consumption to the “carbon intensity” of the country in which the server/client is located. The carbon intensity corresponds to the amount of greenhouse gas emitted per kWh of electricity produced. This information is directly available in some public databases (such as those of the Ember and Energy Institute). In other words, the energy consumption and the energy emissions are perfectly correlated, so we do not provide it.
Indirect communication-related energy consumption
Our experiments exclude the energy consumption indirectly induced by network communications. While it is straightforward to measure the energy consumption of a server or client, estimating the total energy required to transmit information between two devices is more challenging. To the best of our knowledge, no study has yet estimated the complete footprint of client-server communications, as this would require computing the amortized energy consumption of all network devices involved (e.g., routers, switches, antennas, wires) for transmitting a bit of information. Ficher et al. [12] made an initial contribution by studying communication between two servers within the same network backbone, but their work does not provide a generalizable result.
Anyway, the technologies studied in our work induce only minor communication overheads. For example, in HTTPS web browsing, the size of the ciphertext sent to the server is comparable to the plaintext, apart from a one-time secret-key exchange. Thus, their communication overhead is negligible compared to that of “communication-intensive PETs,” such as multiparty computations.
2.3 Is the runtime a good proxy for the energy consumption?
As research papers typically benchmark the runtime of their algorithms, one may want to extrapolate the energy consumption based on such measurements. While runtime is correlated to energy consumption, other parameters influence it. Thus, we cannot trivially deduce the energy consumption based on the runtime.
@@ 2.3 Is the runtime a good proxy for the energy consumption?
This observation shows that specialized energy measurement techniques (such as RAPL [22]) are essential to quantify the consumption and cannot be obtained using simpler methods such as runtime measurement. For example, RAPL-based measurements account for the energy consumption of all cores.
In conclusion, we provide a complementary perspective to existing PET papers because we report energy measurements that cannot be precisely extrapolated from existing experimental results. Our work echoes with recent ML work [19] that promoted a systematic reporting of carbon footprint and energy consumption in ML papers (in addition to existing runtime measurements).
3 Results
This section measures the relative difference between the energy consumption of three PETs and their non-private equivalents. Since it is not feasible to benchmark all PETs in a single study, we focus on three representative technologies selected for their maturity and industrial relevance: FHE, Searchable Encryption, and TLS. On the one hand, these technologies cover different maturity levels: TLS is widely deployed, Searchable Encryption has been actively studied for two decades, and FHE has grown rapidly since the introduction of more efficient schemes [7, 8] less than ten years ago. On the other hand, they each have well-identified applications: encrypted machine learning, encrypted databases, and secure web browsing. With this scope, our goal is both to initiate research on the environmental footprint of PETs and to demonstrate an accessible and reproducible methodology.
This scope naturally leaves room for future studies on other PETs, such as multiparty computation (MPC), functional encryption, or trusted execution environments (TEE). These technologies raise distinct challenges –ranging from massive communication overheads in MPC to indirect hardware-related costs in TEE (see Section 4.1)– that call for dedicated investigation.
@@ 3.1 Experimental setup
As the hardware can slightly change the energy consumption of an algorithm, one may wonder whether we should execute our experiment on different devices. However, remember that we mainly want to measure the relative footprint difference between private and non-private deployments. A standard hardware change would influence the absolute values, but the relative difference should not be significantly changed. This measurement on a single hardware platform aligns with common practices in similar studies on ML algorithms [32, 15].
@@
We use the software CodeCarbon [35] to measure the energy consumption. This software uses a RAPL-based measurement, and measures RAM, CPU, and GPU consumption. We configured CodeCarbon to poll the energy consumption every 1 ms (like in [20]). Our measurement aggregates the computational costs of the client and the server.
Our experiments systematically use the default configuration proposed in the documentation of the tested software. Thus, the measurements reported in our paper correspond to the expected default behavior. For implementation details, we refer to our publicly available codebase:
3.2 Fully Homomorphic Encryption (FHE)
FHE is a type of cryptography scheme [7, 8] allowing to perform arithmetic operations on encrypted data. With the increasing use of ML, there is a growing concern regarding the use of personal data in these systems. To address this problem, FHE has appeared as a nice solution to execute traditional ML pipelines under encryption.
Software libraries. Our goal is to measure the energy consumption of an encrypted ML operations compared to the equivalent plaintext operations. We use the software library Concrete ML developed by the company Zama. This library is stable and promoted by the company for real-world use cases. We compare the encrypted operations provided by Concrete ML to the plaintext ML performed by Scikit Learn (a popular ML framework). Note that Scikit Learn is also used as baseline by Zama in their accuracy benchmarks.
Note that Concrete ML only supports TFHE [8], but there exists another popular FHE scheme in the literature, namely CKKS [7]. Unfortunately, existing CKKS implementations are much more basic than Concrete ML, and do not support high-level operations like ML. While Concrete ML is designed for industrial applications and maintained by a company, other FHE implementations are not as mature. Thus, we only experiment on TFHE, because other implementations either provide insufficient functionalities, or are too immature (i.e., providing an unfair comparison to an industrial software like Concrete ML).
Anyway, our work does not aim to identify the best FHE implementation, or even the best privacy-preserving ML framework in general. We want to estimate an order of magnitude of the energy consumption increase induced by an FHE-based application compared to its plaintext equivalent. Moreover, the FHE literature is particularly active, so reproducibility studies would be needed in the upcoming years to measure the evolution.
@@ 3.2 Fully Homomorphic Encryption (FHE)
3.2.1 Encrypted Machine Learning inference
Among all ML operations, ML inference (i.e., evaluating data on a trained ML model) is a particularly important operation for practitioners. Indeed, many companies have trained a powerful model (e.g., a Large Language Model) and provide “Inference as a Service”: customers send input data and the company returns the model output. To enhance privacy, researchers have designed encrypted ML frameworks [8].
Figure 1: Average energy consumption of encrypted and plaintext ML inference using various classification models (100 samples, 30 features).
Results. Figure 1 presents measurements for various classification and regression models. These results are average measurements over 100 samples; each sample having 30 features. On linear models (e.g., Logistic regression), the encrypted inference is $100\times$ more expensive than the plaintext inference. On tree models (Decision Tree, Random Forest, and XGB), the encrypted inference is at least $100,000\times$ more expensive than the plaintext inference. Encrypted inference on neural networks is also $100,000\times$ more expensive than plaintext.
@@ 3.2.1 Encrypted Machine Learning inference
3.2.2 Encrypted Machine Learning training
Next to inference, ML training is the other operation attracting a lot of privacy concerns.
@@ 3.2.2 Encrypted Machine Learning training
Results. Figure 3(a) compares the encrypted and plaintext training (of a logistic regression) for a varying number of features. The encrypted training is at least $100,000\times$ more expensive and its energy consumption increases faster with the number of features.
Contrary to inference, the logistic regression training is not influenced by the number of samples. Its training algorithm is iterative and process a fixed-sized batch during each iteration. Thus, its cost only depends on the batch size and number of iterations. Figure 3(b) confirms this behavior experimentally.
Discussions. Like encrypted ML inference, secure training induces a massive consumption increase. Similarly, there is room for optimizations.
However, other (non-cryptographic) approaches to privacy-preserving ML training exist. For example, de Reus et al. [32] studied the energy consumption of synthetic data generation in ML context. Synthetic data generation is a pre-processing step used in privacy-preserving statistics and ML: instead of encrypting its private data, the data owner generates synthetic data based on its data and can transmit the synthetic data in plaintext. In this paradigm, the server can then process plaintext data without privacy issues as it has only access to synthetic data (and not the initial private data). Some techniques such as [21] ensure that the synthetic data does not leak private information.
@@
This comparison emphasizes that the PET choice has a major impact on the energy consumption. It also highlights an interesting future work: comparing all privacy-preserving ML paradigms to identify the most energy-efficient one.
3.3 Searchable Encryption
@@
Gui et al. [16] already compared the runtime of SWiSSSE to Redis. We reproduce their benchmark using new metrics: energy.
@@
Results. Figure 4 shows the results of this benchmark for varying database sizes. We observe that the footprint of the encrypted database is nearly $10\times$ higher than the footprint of the plaintext database. Moreover, the footprint of the encrypted database increases slightly faster in function of the database size.
Figure 4: Average energy consumption of encrypted and plaintext queries (SWiSSSE vs. Redis) for varying database sizes (1000 queries).